OpenStudy (dakid88):

The historical walking tour of a town covers a total distance of 5 miles. The map shows that the path of the tour travels 3 inches north, 4 inches east, 2 inches south, and 1 inch west. How many miles does each inch on the map represent? 0.2 miles 0.5 miles 2 miles 10 miles

OpenStudy (dakid88):

@benlindquist

OpenStudy (dakid88):

@jchick @L_N_F_7

OpenStudy (jchick):

Add up the inches

OpenStudy (dakid88):

i did and came out with 15

OpenStudy (mathmale):

Your immediate goal is to find a conversion factor: inches per mile. So, you've come up with 15 inches? And that 15 inches represents how many miles? then what's the conversion factor?

OpenStudy (jchick):

You shouldn't

OpenStudy (jchick):

You should have ten inches and then you divide that by 5

OpenStudy (dakid88):

ok so 10 divided by 5 is 2

OpenStudy (jchick):

So you have your answer

OpenStudy (mathmale):

Yes, except that you need to include units of measurement.

OpenStudy (dakid88):

wht bout u @L_N_F_7

OpenStudy (jchick):

Yes @mathmale

OpenStudy (dakid88):

well i did but came out wit 15

OpenStudy (jchick):

3 + 4 = 7

OpenStudy (anonymous):

u should have 10

OpenStudy (jchick):

7 + 3 = 10

OpenStudy (dakid88):

k i have the answer and the correct answer and it is 0.2

OpenStudy (anonymous):

where u get 7 from @jchick

OpenStudy (dakid88):

lol

OpenStudy (jchick):

The seven is what I had previously added up

OpenStudy (jchick):

Plus the three after

OpenStudy (anonymous):

oh ion know how cause u have to add the 2

OpenStudy (jchick):

(3 + 4) = 7 then add three

OpenStudy (dakid88):

k

OpenStudy (anonymous):

im lost okay i give up

OpenStudy (jchick):

Now what is 10 time 0.5?

OpenStudy (jchick):

10 times 0.5

OpenStudy (jchick):

@dakid88 what is 10 * 0.5?

OpenStudy (dakid88):

5

OpenStudy (dakid88):

k

OpenStudy (dakid88):

can u give me a medal too @jchick

OpenStudy (mathmale):

"How many miles does each inch on the map?" is a RATE (a rate of change). You count 10 inches and know that that 10 inches represents 5 miles. form the RATE OF CHANGE as follows:\[\frac{ 10 inches }{ 5miles }\]

OpenStudy (mathmale):

Next, reduce this. What is 10/5? Write your final result as follows: "Each inch on the map represents ______ miles." or as\[\frac{ ?inches }{ mile }.\]

OpenStudy (jchick):

It needs to be done like this